OpenStudy (anonymous):

How do I find the derivative of y with respect to the appropriate value: y=cos−1(x^2)

OpenStudy (anonymous):

the derivative of y with respect to x?

OpenStudy (anonymous):

yes I believe so

OpenStudy (anonymous):

is the original equation y=cos(-1x^2)

OpenStudy (anonymous):

inverse of cosine so y=cos^1(x^2)

OpenStudy (anonymous):

\[\cos ^{-1}(x ^{2})\] that makes sense now do you know the chain rule?

OpenStudy (anonymous):

not much. Thats why I really need help with this

OpenStudy (anonymous):

the chain rule is a way to find derivatives of composite functions example you could find the derivative of (2x+3)^2 by multiplying out the entire thing and using the power rule or you could use the chain rule the chain rule says that you can substitute in "u" for part of an expresion, take the derivative of the new expression, then multiply it by the derivative of "u" in our example, u=(2x+3) the new equation is u^2 derivative is 2u derivative of u is 2 multiply the derivative of the new equation by the derivative of u and substituting gives you 2(2x+3)(2)=8x+12 if you multiply out the polynomial and use the power rule, you get the same thing try it and see

OpenStudy (anonymous):

so it would be cos^1(2x)?

OpenStudy (anonymous):

you can substitute u=x^2 derivative of u = 2x derivative of cos^-1(u) is -1/(sqrt(1-u^2) if you multiply that by the derivative of u and then re-substitute, you get 2x/(1-sqrt(x^4)) try it and see what you get

OpenStudy (anonymous):

oh okay thank you!

OpenStudy (anonymous):

actually, let me draw out what it should be because my answer actually had another mistake in it

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

do you need more examples?

OpenStudy (anonymous):

well what about for, y=sin^-191-t) ? this one is more complicated.

OpenStudy (anonymous):

sin^-1(191-t)

OpenStudy (anonymous):

i mean y=sin^-1 (1-t) sorry

OpenStudy (anonymous):

what can you set u equal to to make this easier (hint: usually its the argument of the trig function)

OpenStudy (anonymous):

would it be 1?

OpenStudy (anonymous):

how about if we try 1-t

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

because its easier to do sin^-1(u) than it is to do sin^-1(u-t) the u is meant to simplify the equation do you see why we picked 1-t?

OpenStudy (anonymous):

oh okay I see

OpenStudy (anonymous):

what would be the derivative of u=1-t? (hint: its derivative of u with respect to t notationwise, its u'

OpenStudy (anonymous):

is it -1?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

the next step is to find the derivative of inverse sine the formula for d/du sin^-1(u) is 1/sqrt(1-u^2) its similar to inverse cosine, but its positive instead to solve, we take the derivative of the inverse sine, multiply by the derivative of u, then plug in 1-t back in. what do you get?

OpenStudy (anonymous):

is is 1/sqrt (1-1^2)?

OpenStudy (anonymous):